trapezium 66. sq.meter parallal side 28 m and 16 m distance between parallel side

Grade:8

1 Answers

Vijay Mukati
7 years ago
Dear Student,

Although your question is not properly formated, but still i can understand that you are trying to ask the distance between two parallel sides of a trapezium whose area and side length of two parallel sides are given to us.

Since we know that,
Area of trapezium = (1/2)(sum of parallel sides)distance between two parallel sides

Threrefore, the distance between two parallel sides = 2*(area of trapezium)/(a + b)
= 2 * 66/(28 + 16) = 3

Best of Luck!